OpenStudy (anonymous):

A certain amount of money was split between 8 people, and each person got $14.50. How much money was divided?

OpenStudy (anonymous):

it would be 14.50*8

OpenStudy (gorv):

welcome to OS :)

OpenStudy (gorv):

each will have $ 14.50 so total = 14.50+14.50+.......8 times =14.50*8

OpenStudy (anonymous):

Thank you both! :D